Your patient has a confirmed STEMI. You estimate the patient weight at 180 pounds. The medical command physician orders a nitroglycerin drip of 15 mcg/min. You connect a vented microchip set ( 60 gtts/min) to the bottle of Tridil.

Nitroglycerin on hand 25 mg/ 250 mL.

What is the concentration of this bottle of nitroglycerin in mcg?

How many mL/hr are needed?

How many drops per second are needed?
Answer #1

ANSWER.

* Concentration of bottle of nitroglycerin in mcg.

            Nitroglycerin in hand = 25 mg/250 ml.

       1 mg = 1000 mcg

So

     Concentration of nitroglycerin in mcg = 25 * 1000

                                                                       = 25000 mcg.

* Calculation of ml/hr.

           ml/hr = Volume(ml)/Time(hr).

   Calculation of volume(ml).

       Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

Here

     Desired dose(D) = 15 mcg/min.

     Dose in hand(H) = 25000 mcg.

          Quantity(Q) = 250 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                   = 15/25000 * 250

                                                   = 0.0006 * 250

                                                   = 0.15

                Time = 1 min

   Convert ml/min to ml/hr by multiply with 60

          ml/hr = 0.15 * 60

                         = 9 ml/hr.

Calculation of drops per second.

        Drops per min = Volume(ml)/Time(min) * Drop factor(gtt/ml)

Here

        Volume(ml) = 9 ml

        Time = 60 min (1 hr).

       Drop factor = 60 gtt/ml.

As per above formula

           Drops per min = 9/60 * 60

                                   = 0.15 * 60

                                   = 9 ml.

Convert drop per min to drop per sec by divide with 60

       So

              Drops per sec = 9/60

                                      = 0.15 drops per sec.